OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = (x-8)/(x+7) g(x)= (-7x-8)/(x-1)

OpenStudy (anonymous):

@bridgetx516x @timo86m

OpenStudy (anonymous):

@abb0t @beketso @Chooch146 @drummerboyrevived @emily2013 @farmergirl411 @Giovanni @help123please. @Meepi

OpenStudy (anonymous):

@Mertsj @Milani123

OpenStudy (anonymous):

Someone PLz Help

OpenStudy (anonymous):

first f(g(x)) = f((-7x-8)/(x-1)) meaning where u see x in the expression of f u replace it with ((-7x-8)/(x-1))

OpenStudy (anonymous):

i already found f(g(x)) i cant find g(f(x))

OpenStudy (anonymous):

ok.give a minute

OpenStudy (anonymous):

@Chooch146 help plz

OpenStudy (anonymous):

@zepdrix

OpenStudy (mertsj):

Can you find the f (g(x))?

OpenStudy (anonymous):

i already have f((-7x-8/x-1) = [ (-7x-8)/(x-1) - 8 ] / [ (-7x-8)/(x-1) + 7 ] = -15x/-15 = x

OpenStudy (mertsj):

\[f(g(x))=\frac{\frac{-7x-8}{x-1}-8}{\frac{-7x-8}{x-1}+7}\times \frac{x-1}{x-1}\frac{-7x-8-8(x-1)}{-7x-8+7(x-1)}\]

OpenStudy (mertsj):

= \[\frac{-7x-8-8x+8}{-7x-8+7x-7}=\frac{-15x}{-15}=x\]

OpenStudy (mertsj):

So we have shown that the f(g(x))=x. Now we must show that the g(f(x))=x

OpenStudy (anonymous):

ok. thats is the one i can't show

OpenStudy (anonymous):

@Mertsj could u help me show g(f(x)) PLZ???

OpenStudy (mertsj):

\[g(f(x))=\frac{-7\times\frac{x-8}{x+7}-8}{\frac{x-8}{x+7}-1}\times\frac{x+7}{x+7}=\frac{-7(x-8)-8(x+7)}{x-8-1(x+7)}\]

OpenStudy (mertsj):

Can you finish it now?

OpenStudy (anonymous):

would the answer be 15? @Mertsj

OpenStudy (mertsj):

I believe that your instructions say that the g(f(x))=x

OpenStudy (anonymous):

then i must have finished it wrong. where did i mess up

OpenStudy (mertsj):

Post what you did and I will see.

OpenStudy (anonymous):

x-8 cancels out and x-7 cancels out which gives you -7-8/-1 which leaves you with 15

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

x-8 does not cancel out. Only factors can be cancelled.

OpenStudy (mertsj):

What is -7(x-8)

OpenStudy (anonymous):

-7x+56 @Mertsj

OpenStudy (mertsj):

What is -8(x+7)?

OpenStudy (anonymous):

-8x -56 @Mertsj

OpenStudy (mertsj):

So the numerator: -7(x-8)-8(x+7)=-7x+56-8x-56

OpenStudy (mertsj):

Simplify that some more.

OpenStudy (anonymous):

-15x

OpenStudy (mertsj):

So the numerator is -15x. Now do the denominator

OpenStudy (anonymous):

denominator is 15

OpenStudy (mertsj):

Try again

OpenStudy (anonymous):

than k u for walking me thru this!!!!!

OpenStudy (anonymous):

-15 sorry

OpenStudy (mertsj):

What is the denominator?

OpenStudy (mertsj):

So now we have: \[\frac{-15x}{-15}\]

OpenStudy (mertsj):

What does that equal?

OpenStudy (anonymous):

x

OpenStudy (mertsj):

So now we have shown that f(g(x))=x And g(f(x))=x And that is the definition of inverse functions so we have shown that f and g are indeed inverse functions.

OpenStudy (anonymous):

THANK U

OpenStudy (mertsj):

yw